Use pair data values and the summations given to
x:& 1 & & & & 3 & 5 & & 7 & & 9
y:& 10& & & 8& & 7& & & 4& & & 2

Σx= 25;& & & Σy= 31;& & & & & Σxy= 115:& & & & Σx² = 165

a. Find the equation for the least squares regression line.
b. Predict the value of y when x is 6 using your equation
1
Expert's answer
2018-07-30T09:24:07-0400
1. The average x and y are
x = 25/5 = 5
y = 31/5 = 6.2
2. The regression coefficient is


thus the equation of the least squares would be

https://latex.codecogs.com/gif.latex?(x-5) = -1(y - 6.2) \ y = -x + 11.2
When x= 6,
y = -6 + 11.2 = 5.2